# Consumption Smoothing in Metropolis: Evidence from Working-class Households in Prewar Tokyo

### v12

@@ B.12 Risk Preferences
Schulhofer-Wohl (2011) shows a specific case (utility function) that the heterogeneous risk preference may lead to bias in the estimation of the income elasticity under the standard risk-sharing regressions and proposes a parametric test for that heterogeneity. This test is valid under a particular utility function. Specifically, the bias could be raised in the case that the coefficient of relative risk aversion interacts with the aggregate shock in the first-order condition for consumption because the interaction means that the aggregate shock cannot be removed using the time-fixed effect in the risk-sharing regressions.104104104Specifically, Schulhofer-Wohl (2011, pp. 928–929) considers the utility maximization problem under the following weighted sum of discounted expected utilities: $\sum_{i}\dot{\alpha}_{i}E\left[\sum_{t}\dot{\beta}^{t}\frac{[c_{i,t}(s_{t})]^{1-\dot{\gamma}_{i}}}{1-\dot{\gamma}_{i}}\right],$ (17) where $\dot{\alpha}_{i}$ is Parate weight and $\dot{\gamma}_{i}$ is the coefficient of relative risk aversion for household $i$. The log-transformed first-order condition becomes: $\text{log}c_{i,t}=\frac{\text{log}\dot{\alpha}_{i}}{\dot{\gamma}_{i}}-\frac{\text{log}\dot{\lambda}_{t}}{\dot{\gamma}_{i}},$ (18) where $\dot{\lambda}_{t}$ indicates the Lagrangian multiplier. This result implies that the effects of the aggregate shocks ($\dot{\lambda}_{t}$) depend on the household’s attitudes toward risk and that the first-differencing does not remove the second term in equation 18. As explained, this setting differs from the one used in this current study. Mozzoco and Saini (2012) consider a more general case for testing the homogeneity of risk preferences. Their test requires a panel dataset with a very long time dimension that is not applicable to my THBS dataset.105105105The data used in their study covers the household-level monthly panel dataset surveyed in rural India between 1975 and 1985. The households with fewer than 80 data points are not included in the homogeneity tests (Mozzoco and Saini 2012, p. 453). See also Shrinivas and Fafchamps (2018) and Mozzoco and Saini (2018) for the subsequent discussions.
@@
An alternative way to test potential heterogeneity in the risk preference is to use direct measures on the risk preferences of workers. Schulhofer-Wohl (2011) uses the results of experimental risk-tolerance questions on the workers’ risk preferences in the Health and Retirement Study to test whether the correlations between earnings and aggregate shocks differ across high- and low-risk-tolerance workers.106106106This question is asking if the respondents would be willing to take a risk to get a new job with a higher income, even if there was a 50-50 chance of a reduction in income by a third (Schulhofer-Wohl 2011, P. 931). They found that the high-risk-tolerance workers have more procyclical incomes than the low-risk-tolerance workers.
@@
Since the THBS has no experimental questions, it is impossible to apply the same test. Given this limitation—that the direct test using experimental risk-tolerance question is unavailable—I compare the correlations between the heads’ income and aggregate consumption over sub-groups to obtain insights into the potential difference in the heads’ attitudes toward risks. For household $i$ and semi-month $t$, the specification is characterized as follows:
@@
$d_{i}$ is an indicator variable that takes one if a household $i$ belongs to a specific group that might have had different preferences on their jobs than the other heads in the sample. Firstly, I test the hypothesis that the heads who work in the large-scale factories have different risk preferences. Although there are no specific anecdotes behind this, the workers in the large-scale factories might have had a lower risk tolerance than those in the small- and medium-scale enterprises. Thirteen heads who worked in the Ishikawajima shipyard and governmental factory are classified into this group. Secondly, I test the hypothesis that the heads who work in the smithing and the non-machinery factories have different risk preferences. As described in Online Appendix A.5, skilled workers in the smithing sector could be small business owners.107107107Workers in the canning factories might have had a different preference on the risk preferences (Online Appendix A.6). However, my THBS sample does not include the workers in canning factories. The workers in non-machinery factories might also have different job preferences than the machinery factory workers. This group has only three heads: one worked in a smithing factory, and two worked in food and paper-making factories.
This finding is consistent with the fact that these heads have similar characteristics because the THBS focused on skilled factory workers in a representative manufacturing area, Tsukishima. Therefore, it is plausible that the THBS households had similar attitudes toward the risks.
@@
Notes: This table shows the results of equation 19 for the semi-monthly panel dataset. ‘Aggregate Consumption’ indicates the cross-sectional sample average of the total expenditures. ‘Large Scale Factory’ indicates an indicator variable that takes one if the head worked in the Ishikawajima Shipyard or governmental factory. ‘Smithing/Non-machinery’ indicates an indicator variable that takes one if the head worked in the smithing, food, or paper-making factories. All the regressions include the family size variable and household fixed effect. Columns 1–2 show the results for the entire sample period. Columns 3–4 show the results for the period between January and September 1919 to remove year-month cells with the smaller (cross-sectional) number of households (Online Appendix Figure B.4).
@@ C.1 Conceptual Framework
Consider the economy where household $i=1,...,N$ receives an uncertain income at time $t$ depending on the state of the world $s_{t}\in S_{t}$ as $y_{i,t}(s_{t})$. For infinite time horizon, the state $s_{t}$ realizes with probability $\pi(s_{t})\in[0,1]$. The weighted sum of the expected lifetime utility of $N$ households in the economy is expressed as:
@@
where $\omega_{i}$ is the social planner’s weight representing the reciprocal of the marginal utility of each household ($10<\omega_{i}<$ 1 and $\sum_{i=1}^{N}\omega_{i}=1$), $10<\rho^{t}<$ 1 is the discount factor, $c_{i,t}(s_{t})$ is consumption, and $\theta_{i,t}(s_{t})$ is a preference shock. For all states $s_{t}$, the aggregate resource constraint is given by:
@@
where $c_{i,t}$ is consumption, $y_{i,t}$ is disposable income, $x_{i,t}$ is family size, $\mu_{i}$ is the household fixed effect, $\phi_{t}$ is the time fixed effect, and $u_{i,t}$ is a random error term. The estimate of $\gamma$ shall range from zero (i.e., idiosyncratic shocks are perfectly insured) to one (for the absence of insurance).
@@
where $\tilde{y}_{i,t}$ is the head’s earning, $x_{i,t}$ is the family size control, $\nu_{i}$ indicates the household-fixed effect, $\zeta_{t}$ indicates the time-fixed effect, and $\epsilon_{i,t}$ is a random error term. I herein use the head’s earnings ($\tilde{y}_{i,t}$) as the idiosyncratic shock variable instead of the disposable income.108108108In the analysis, I consider the labor-supply adjustment variable and sales of miscellaneous assets as the additional dependent variables in addition to the net income variables $r_{i,t}$ (Section 4.2). The former variable is the total earnings of the family members except for the head, meaning that the simultaneous issue arises if I use the disposable income in this equation. The regression is separately run for each component in the net income vector ($\bm{z}$).
@@ C.2 Raw Relationship between Income and Expenditures
Figure C.2 summarizes the scatter plots between the LD of the semi-monthly disposable income and expenditures. Figure C.2a illustrates the relationship between the LD of the semi-monthly disposable income and total expenditure. Figures C.2bC.2l illustrate the correlations between the LD of the semi-monthly disposable income and 11 expenditure subcategories.
Given this technical issue, the adjusted monthly panel dataset is suitable for estimating the income elasticity for the rent category. In fact, Figure C.4c confirms that the scatter plot from the adjusted monthly variation shows a flatter relationship.
(a) Total expenditure
@@
Notes: The distribution of the log-difference in semi-monthly total expenditure is shown in Figure C.1a. Figures C.1bC.1l show the distributions of the log-differences in semi-monthly expenditure for the 11 subcategories listed in panel A of Table 3.
Source: Created by the author using the THBS sample.
(a) Total expenditure
@@
Notes: Figure C.2a illustrates the relationship between the log-difference in semi-monthly disposable income and expenditure. Figures C.2bC.2l illustrate the relationships between the log-differences in semi-monthly disposable income and the 11 expenditure subcategories listed in Panel A in Table 3. The range of the y-axis is fixed across all the figures for comparability.
Source: Created by the author using the THBS sample.
(a) Total expenditure
@@
Notes: The distribution of the log-difference in adjusted monthly total expenditure is shown in Figure C.3a. Figures C.3bC.3l show the distributions of the log-differences in adjusted monthly expenditure for the 11 subcategories listed in panel A of Table 3.
Source: Created by the author using the THBS sample.
(a) Total expenditure
@@
Notes: Figure C.4a illustrates the relationship between the log-difference in semi-monthly disposable income and expenditure. Figures C.4bC.4l illustrate the relationships between the log-differences in semi-monthly disposable income and the 11 expenditure subcategories listed in Panel A in Table 3. The range of the y-axis is fixed across all the figures for comparability.
Sources: Created by the author using the THBS sample.
C.3 Retail Sales in Tokyo City: A Retail Store Survey of 1935

### v14

@@ B.12 Risk Preferences
Schulhofer-Wohl (2011) shows a specific case (utility function) that the heterogeneous risk preference may lead to bias in the estimation of the income elasticity under the standard risk-sharing regressions and proposes a parametric test for that heterogeneity. This test is valid under a particular utility function. Specifically, the bias could be raised in the case that the coefficient of relative risk aversion interacts with the aggregate shock in the first-order condition for consumption because the interaction means that the aggregate shock cannot be removed using the time-fixed effect in the risk-sharing regressions. [^87] Specifically, Schulhofer-Wohl (2011, pp. 928–929) considers the utility maximization problem under the following weighted sum of discounted expected utilities: $\sum_{i}\dot{\alpha}_{i}E\left[\sum_{t}\dot{\beta}^{t}\frac{[c_{i,t}(s_{t})]^{1-\dot{\gamma}_{i}}}{1-\dot{\gamma}_{i}}\right],$ (17) where $\dot{\alpha}_{i}$ is Parate weight and $\dot{\gamma}_{i}$ is the coefficient of relative risk aversion for household $i$. The log-transformed first-order condition becomes: $\text{log}c_{i,t}=\frac{\text{log}\dot{\alpha}_{i}}{\dot{\gamma}_{i}}-\frac{\text{log}\dot{\lambda}_{t}}{\dot{\gamma}_{i}},$ (18) where $\dot{\lambda}_{t}$ indicates the Lagrangian multiplier. This result implies that the effects of the aggregate shocks ($\dot{\lambda}_{t}$) depend on the household’s attitudes toward risk and that the first-differencing does not remove the second term in equation 18. As explained, this setting differs from the one used in this current study. Mozzoco and Saini (2012) consider a more general case for testing the homogeneity of risk preferences. Their test requires a panel dataset with a very long time dimension that is not applicable to my THBS dataset. [^88] The data used in their study covers the household-level monthly panel dataset surveyed in rural India between 1975 and 1985. The households with fewer than 80 data points are not included in the homogeneity tests (Mozzoco and Saini 2012, p. 453). See also Shrinivas and Fafchamps (2018) and Mozzoco and Saini (2018) for the subsequent discussions.
@@
An alternative way to test potential heterogeneity in the risk preference is to use direct measures on the risk preferences of workers. Schulhofer-Wohl (2011) uses the results of experimental risk-tolerance questions on the workers’ risk preferences in the Health and Retirement Study to test whether the correlations between earnings and aggregate shocks differ across high- and low-risk-tolerance workers. [^89] This question is asking if the respondents would be willing to take a risk to get a new job with a higher income, even if there was a 50-50 chance of a reduction in income by a third (Schulhofer-Wohl 2011, P. 931). They found that the high-risk-tolerance workers have more procyclical incomes than the low-risk-tolerance workers.
@@
Since the THBS has no experimental questions, it is impossible to apply the same test. Given this limitation, I compare the correlations between the heads’ income and aggregate consumption over sub-groups to obtain insights into the potential difference in the heads’ attitudes toward risks. For household $i$ and semi-month $t$, the specification is characterized as follows:
@@
$d_{i}$ is an indicator variable that takes one if a household $i$ belongs to a specific group that might have had different preferences on their jobs than the other heads in the sample. Firstly, I test the hypothesis that the heads who work in the large-scale factories have different risk preferences. Although there are no specific anecdotes behind this, the workers in the large-scale factories might have had a lower risk tolerance than those in the small- and medium-scale enterprises. Thirteen heads who worked in the Ishikawajima Shipyard and governmental factory are classified into this group. Secondly, I test the hypothesis that the heads who work in the smithing and the non-machinery factories have different risk preferences. As described in Online Appendix A.5, skilled workers in the smithing sector could be small business owners. [^90] Workers in the canning factories might have had a different preference on the risk preferences (Online Appendix A.6). However, my THBS sample does not include the workers in canning factories. The workers in non-machinery factories might also have different job preferences than the machinery factory workers. This group has only three heads: one worked in a smithing factory, and two worked in food and paper-making factories.
@@
Notes: This table shows the results of equation 19 for the adjusted semi-monthly panel dataset. ‘Aggregate Consumption’ indicates the cross-sectional sample average of the total expenditures. ‘Large Scale Factory’ indicates an indicator variable that takes one if the head worked in the Ishikawajima Shipyard or governmental factory. ‘Smithing/Non-machinery’ indicates an indicator variable that takes one if the head worked in the smithing, food, or paper-making factories. All the regressions include the family size variable and household fixed effect. Columns 1–2 show the results for the entire sample period. Columns 3–4 show the results for the period between January and September 1919 to remove year-month cells with the smaller (cross-sectional) number of households (Online Appendix Figure B.4).
This finding aligns with the similarity of these household heads, as the THBS targeted skilled factory workers within a geographically narrow area. Therefore, it is plausible that the THBS households had similar attitudes toward the risks.
@@ C.1 Conceptual Framework
Consider the economy where household i=1,,Ni=1,...,N receives an uncertain income at time $t$ depending on the state of the world $s_{t}\in S_{t}$ as $y_{i,t}(s_{t})$. For infinite time horizon, the state $s_{t}$ realizes with probability $\pi(s_{t})\in[0,1]$. The weighted sum of the expected lifetime utility of $N$ households in the economy is expressed as:
@@
where $\omega_{i}$ is the social planner’s weight representing the reciprocal of the marginal utility of each household ($0<\omega_{i}<1$ and $\sum_{i=1}^{N}\omega_{i}=1$), $0<\rho^{t}<1$ is the discount factor, $c_{i,t}(s_{t})$ is consumption, and $\theta_{i,t}(s_{t})$ is a preference shock. For all states $s_{t}$, the aggregate resource constraint is given by:
@@
where $c_{i,t}$ is consumption, $y_{i,t}$ is disposable income, $x_{i,t}$ is family size, $\mu_{i}$ is the household fixed effect, $\phi_{t}$ is the time fixed effect, and $u_{i,t}$ is a random error term. The estimate of $\gamma$ shall range from zero (i.e., idiosyncratic shocks are perfectly insured) to one (for the absence of insurance). [^91] I use the conventional within estimator. The bias-corrected estimator for the two-way fixed-effects model in a difference-in-differences (TWFE-DID) design with continuous treatment over multiple periods remains underdeveloped (de Chaisemartin and D’Haultfœuille 2023). Therefore, previous studies have not yet incorporated recent theoretical developments in the TWFE-DID framework.
@@
where $\tilde{y}_{i,t}$ is the head’s earning, $x_{i,t}$ is the family size control, $\nu_{i}$ indicates the household-fixed effect, $\zeta_{t}$ indicates the time-fixed effect, and $\epsilon_{i,t}$ is a random error term. I herein use the head’s earnings ($\tilde{y}_{i,t}$) as the idiosyncratic shock variable instead of the disposable income. [^92] I consider the labor-supply adjustment variable and sales of miscellaneous assets as the additional dependent variables in addition to the net income variables $r_{i,t}$ (Section 4.2). The former variable is the total earnings of the family members except for the head, meaning that the simultaneous issue arises if I use the disposable income in this equation. The regression is separately run for each component in the net income vector ($\bm{z}$).
@@ C.2 Raw Relationship between Income and Expenditures
(a) Total
@@
Notes: The distribution of the log-difference in adjusted semi-monthly total expenditure is shown in Figure C.1a. Figures C.1bC.1l show the distributions of the log-differences in adjusted semi-monthly expenditure for the 11 subcategories listed in panel A of Table 3. Source: Created by the author using the THBS sample.
(a) Total
@@
Notes: Figure C.2a illustrates the relationship between the log-difference in adjusted semi-monthly disposable income and expenditure. Figures C.2bC.2l illustrate the relationships between the log-differences in adjusted semi-monthly disposable income and the 11 expenditure subcategories listed in Panel A in Table 3. The range of the y-axis is fixed across all the figures for comparability. Source: Created by the author using the THBS sample.
(a) Total
@@
Notes: The distribution of the log-difference in adjusted monthly total expenditure is shown in Figure C.3a. Figures C.3bC.3l show the distributions of the log-differences in adjusted monthly expenditure for the 11 subcategories listed in panel A of Table 3. Source: Created by the author using the THBS sample.
(a) Total
@@
Notes: Figure C.4a illustrates the relationship between the log-difference in monthly disposable income and expenditure. Figures C.4bC.4l illustrate the relationships between the log-differences in monthly disposable income and the 11 expenditure subcategories listed in Panel A in Table 3. The range of the y-axis is fixed across all the figures for comparability. Sources: Created by the author using the THBS sample.
Figure C.2 summarizes the scatter plots between the LD of the adjusted semi-monthly disposable income and expenditures. Figure C.2a illustrates the relationship between the LD of the adjusted semi-monthly disposable income and total expenditure. Figures C.2bC.2l illustrate the correlations between the LD of the adjusted semi-monthly disposable income and 11 expenditure subcategories.
Given this technical issue, the monthly panel dataset is suitable for estimating the income elasticity for the rent category. In fact, Figure C.4c confirms that the scatter plot from the monthly variation shows a flatter relationship.
C.3 Testing Heterogeneity: Accessibility to Retailers
The approximate locations of retail shops for daily necessities are reported in the official document and shown in Figure C.5. Density varies across lots: the number of retail shops per $1,000$ population is $8.8$ in the Tsukuda area, $7.8$ in Lot 1, and $4.6$ in Lot 2. Lot 2 has a lower density because it is relatively newer than the others (Online Appendix A.3).
First, I include the density of retail shops in each lot as a control variable in Equation 2. Because this variable is time-constant, I interact it with the time dummies. Column 1 of Table C.1 shows an estimate of 0.132-0.132, which is close to the baseline estimate of 0.128-0.128 (Table C.7).
Second, I expand Equation 2 by adding an interaction term between the head’s earnings and retail shop density. In this expanded model, a statistically significantly negative coefficient on the interaction term would indicate that households in denser areas rely more on credit purchases. However, Column 2 shows that the estimated coefficient on the interaction term is 0.004-0.004 and not statistically significant. In Column 3, I use an alternative definition of density, adopting the number of households as the denominator to account for variation in family size. The estimated coefficient on the interaction term remains near zero (0.002-0.002). In both columns, the marginal effect of the head’s earnings is estimated at 0.137-0.137, similar to the estimate in Column 1.
The total length of the residential area in Tsukishima is roughly $1.4$ kilometers with a width of $500$ meters (Chūō Ward 1994, p. 38). Thus, although the living areas are divided into three lots, residents can easily access shops in different lots on foot (Department of Health, Ministry of the Interior 1923a, pp. 40–46).
Figure C.5: Retail Shops for Daily Necessities: Food and Fuel Retailers
Notes: This map shows the locations of retail shops for daily necessities (seikatsu hitsujyuhin hanbaiten) in Tsukishima. Cross mark: rice shop (kome ya). Circle with a cross: liquor shop (saka ya). Black triangle: greengrocer (yao ya) Circle with a minus: fish store (sakana ya). Black square: tōfu store. Half-black circle: firewood and charcoal store (shintan ya). Square with a cross: butcher store (niku ya). White circle: sugar, pickles, and dried food store (satō, tsukemono, kanbutsu ya). Black circle: tsukudani shop. tsukudani is a food simmered in sweet soy sauce. Half circle: tea shop (cha ya). White triangle: goshikiage shop. Note: Lot 3, the southernmost lot, had no households in 1919. Goshikiage ya is a fried fish paste containing vegetables. Source: Department of Health, Ministry of the Interior 1923c, fourth map. The tone was adjusted by the author using Adobe Photoshop 26.6.1.
Table C.1: Testing Heterogeneity: Density of the Retailers in Each Lot
Dependent Variable: Net Credit Purchase (income minus expenditure, yen)
Semi-monthly Panel Adjusted Monthly Panel
(1) (2) (3) (4) (5) (6)
Head’s Earnings 0.132-0.132*** 0.107-0.107 0.099-0.099 0.026-0.026 $0.010$ $0.013$
$[0.044]$ $[0.117]$ $[0.135]$ $[0.061]$ $[0.160]$ $[0.190]$
Head’s Earnings $\times$ Density 0.004-0.004 0.002-0.002 0.003-0.003 0.001-0.001
$[0.020]$ $[0.007]$ $[0.026]$ $[0.009]$
Marginal Effect ($\partial r_{it}/\partial\tilde{y}_{it}$) 0.137-0.137*** 0.137-0.137*** 0.008-0.008 0.008-0.008
$[0.045]$ $[0.044]$ $[0.051]$ $[0.051]$
Density $\times$ Time FEs Yes No No Yes No No
Family Size Yes Yes Yes Yes Yes Yes
Denominator of the Density Population Population Households Population Population Households
Observations 279 279 279 120 120 120
Notes: Columns 1 and 4 report the estimated coefficient on the head’s earnings in the expanded regression of Equation 2, which includes interaction terms between retailer density and time dummies as additional controls. Columns 2, 3, 5, and 6 present results from the expanded regression of Equation 2, which includes the interaction between the head’s earnings and retailer density. The density variable in Columns 2 and 5 is measured per population, whereas it is measured per household in Columns 3 and 6. Columns 1–3 show results for the semi-monthly panel, while Columns 4–6 show results for the adjusted monthly panel. “Marginal Effect ($\partial r_{it}/\partial\tilde{y}_{it}$)” indicates the marginal effect of the head’s earnings at the mean of the density variable. All regressions control for family size. Households with missing residence information are excluded from the analytical sample. Results remain unchanged if these households are assumed to reside in any of the three lots. A few households living on the mainland in Kyobashi Ward are assigned to the nearest lot of Tsukishima. Sources: See the main text for the THBS sample. Data on the number of people and households in the three lots are from the Department of Health, Ministry of the Interior (1923b, p. 2). The number of retail shops in each lot is from the Department of Health, Ministry of the Interior (1923c, fourth map).
C.4 Testing the Roles of Cooperatives and Pawnshops
Table C.2: Testing Roles of the Cooperatives and Pawnshops
Panel A: Estimates from the Baseline Specification
Dependent Variable:
Prices of Purchased Items
from Cooperatives (yen) Income from Pawnshops (yen)
(1) (2) (3) (4)
Head’s Earnings 0.001-0.001 0.006-0.006 $0.000$ $0.005$
$[0.002]$ $[0.004]$ $[0.002]$ $[0.006]$
Type of Panels Semi-monthly Monthly Semi-monthly Monthly
Family Size Yes Yes Yes Yes
Observations 289 124 289 120
Panel B: Testing Heterogeneity using Density of the Pawnshop in Each Lot
Dependent Variable: Income from Pawnshops (yen)
Semi-monthly Panel Adjusted Monthly Panel
(1) (2) (3) (4) (5) (6)
Head’s Earnings 0.003-0.003 $0.013$ $0.018$ $0.001$ $0.113$ $0.101$
$[0.003]$ $[0.022]$ $[0.025]$ $[0.004]$ $[0.126]$ $[0.115]$
Head’s Earnings $\times$ Density 0.079-0.079 0.032-0.032 0.686-0.686 0.182-0.182
$[0.129]$ $[0.044]$ $[0.753]$ $[0.203]$
Marginal Effect ($\partial r_{it}/\partial\tilde{y}_{it}$) $0.001$ $0.001$ $0.006$ $0.007$
$[0.002]$ $[0.003]$ $[0.009]$ $[0.010]$
Density $\times$ Time FEs Yes No No Yes No No
Family Size Yes Yes Yes Yes Yes Yes
Denominator of the Density Population Population Households Population Population Households
Observations 279 279 279 120 120 120
Notes: Panel A: This panel reports results from Equation 2, using prices of purchased items from cooperatives (Columns 1–2) or temporary income from pawnshops (Columns 3–4) as the dependent variable. Columns 1 and 3 use the semi-monthly panel, while Columns 2 and 4 use the adjusted monthly panel. All regressions control for family size. Panel B: Columns 1 and 4 present the estimated coefficient on the head’s earnings in the expanded regression of Equation 2, which includes interaction terms between pawnshop density and time dummies as additional controls. Columns 2, 3, 5, and 6 present results from the expanded regression, accounting for the interaction between the head’s earnings and pawnshop density. The density variable in Columns 2 and 5 is measured per population, while in Columns 3 and 6, it is measured per household. Columns 1–3 show results for the semi-monthly panel, and Columns 4–6 show results for the adjusted monthly panel. “Marginal Effect ($\partial r_{it}/\partial\tilde{y}_{it}$)” indicates the marginal effect of the head’s earnings at the mean of the density variable. Households with no address are excluded from the analytical sample. Results are unchanged if these households are assumed to reside in any of the three lots. A few households living on the mainland in Kyobashi Ward are assigned to the nearest lot of Tsukishima. Sources: See the main text for the THBS sample. Data on the number of people and households in the three lots are from the Department of Health, Ministry of the Interior (1923b, p. 2). The number of pawnshops in each lot is from the Department of Health, Ministry of the Interior (1923c, 15th map).
In the THBS sample, five households used cooperatives to buy daily commodities in a few semi-monthly cells. Although this is a small portion, cooperatives may have offered some compensation to their members. To examine this possibility, I use the total prices of items purchased from cooperatives as the dependent variable in Equation 2. If such purchases increase following an idiosyncratic shock, the estimated coefficient on the head’s income would turn negative and become statistically significant. Columns 1 and 2 in Panel A of Table C.2 show that the estimates are negative but close to zero and statistically insignificant in both the semi-monthly and adjusted monthly panels.
Similarly, a few THBS households used pawnshops in a couple of semi-monthly cells. I examine the role of pawnshops using the nominal income from pawnshops in Equation 2.939393Nominal income is used instead of net income because the THBS budget book summarizes payments in a column labeled payment (bensai), which includes amounts paid not only to pawnshops but also to other lending institutions. If workers rely on pawnshops to mitigate hardships, the estimated coefficient on the head’s income would be negative. Columns 3 and 4 in Panel B show that the estimates are near zero and statistically insignificant in both the semi-monthly and adjusted monthly panels. This result aligns with the main finding for the borrowing category (Table 5).
In Panel B of the same table, I examine potential heterogeneity in pawnshop usage following the approach in Table C.1. The official report provides the locations of pawnshops, allowing me to calculate their density, measured as the number of pawnshops per $1,000$ people, in each lot: $0.13$, $0.18$, and $0.16$ in the Tsukuda area, Lot 1, and Lot 2, respectively. Column 1 includes density as a control variable in Equation 2. Column 2 adds an interaction term between the head’s income and the density variable. Column 3 tests sensitivity by using the number of households as an alternative denominator for density. In all columns, the estimated coefficients are near zero and not statistically significant. Columns 4–6, using the adjusted monthly panel, produce the same results. This pattern aligns with the findings in Table C.1, consistent with Tsukishima’s geographically narrow layout.
C.5 Retail Sales in Tokyo City: A Retail Store Survey of 1935